QUESTION 1:

You are testing H0: µ = 100 against Ha: µ < 100 based on an SRS of 18 observations from a Normal population. The data give x¯¯¯x¯ = 8.3 and s = 5.


The value of the t statistic (±0.01) is   

QUESTION 2:

You have an SRS of 14 observations from a Normally distributed population.


What critical value (±±0.001) would you use to obtain a 99.5% confidence interval for the mean μμ of the population?    
Answer #1

Answer:-- Date:--24/07/2019 9; -.7-4 – 8.3 – 100 = -77.810

2). n = 14 Degrees of freedom = df=n - 1 = 14 - 1 = 13 At 99.5% confidence level the tis, a=1-99.5%= 1-0.995 = 0.005 al 2 = 0